### v1

@@ 3.1 Data
The primary data used to construct the election shocks come from Iowa Electronic Markets (IEM), a futures contract market operated by the University of Iowa for research purposes. IEM publicly provides data on its winner take all futures betting markets for presidential races. Betting markets have been shown to more accurately predict elections than polling data (Berg et al., 2008). Relative to other betting markets, IEM has several desirable properties: (i) bets are capped at \$500, mitigating the influence of single traders; (ii) the markets have consistently high trading volumes, especially in the last few months of an election; (iii) the data are at a daily frequency, which is vital for studying stock markets; and (iv) market data is available for every presidential election since 1988999I focus on elections starting in 2000 due to the lack of availability of trading volumes and sectoral price indices for earlier elections., and in all election years data are available starting in June or earlier and extend through election day in November. One drawback of these data are that the payouts are based on which party wins the popular vote rather than the electoral college. In other words, the implied probability of a candidate winning based on these futures prices are biased if her chances of winning the popular vote are different from her chances of winning the election. Indeed, this is particularly relevant for recent elections, e.g., based on other betting markets Harris’s predicted probability of winning the popular vote was much higher than her probability of winning. [^10] For example, odds from Polymarket.com the day before the election implied Harris had a 71% chance to win the popular vote, but a 41% chance to win the presidency. While the last three elections had electoral college bias in favor of the Republican candidates, there is no clear historical systematic bias (Erikson et al. 2020; Coleman and Kondik, 2023). With this in mind, to construct the election shocks I rely on the implicit correlation between innovations to both a candidates’ probability of winning the popular vote and the probability that they win electoral college. Aside from adding additional noise, this does not present any clear threats to identification. However, where relevant I nuance the discussion to account for this caveat.
@@
Figure 4 plots the 2024 Republican and Democratic candidates’ implied winning probabilities from June 1st through November 4th, the day before the election. The probabilities sum to 100, i.e., an increase to the Republicans’ probabilities are associated with a 1-to-1 decrease in the Democrats’ probabilities. Democratic candidates were each favorites to win the popular vote throughout the sample, although there is significant heterogeneity in the magnitude of this advantage. Vertical lines indicate notable events that are associated with substantial movements in the data. The first and most notable event is the aforementioned debate between Biden and Trump, which is associated with a contemporaneous 15pp increase to Trump’s probability of winning, albeit leveling off roughly 10pp higher than pre-debate levels. On July 11, Biden held a news conference that received mixed reactions, but overall was seen as a positive signal relative to his debate performance. This event is associated with a moderate increase to the Democrats’ odds of presidency. Just two days later, Trump survived an assassination attempt during a rally in Pennsylvania, which bolstered his election probabilities. A medium-term increase in the Democrats likelihoods appears to be linked to Biden dropping out of the race and endorsing Harris on July 21. Finally, the aforementioned debate between Harris and Trump is associated with a steady increase in the Democrats’ odds, which peak about 16pp higher roughly two weeks after the debate. See Figure B1 for the probabilities of other elections in the sample and Table C2 for information on other identified events. Overall, these data appear to adequately track fluctuations in each party’s election probabilities around significant events.
@@ 3.2 Construction of the Election Shocks
The goal of this section is to construct a shock series that captures the types of fluctuations in election probabilities that are emphasized in Figure 4. However, a significant challenge to this exercise is the possibility that movements in election probabilities also reflect changes in voter preferences given changes to the state of financial markets or the real economy. For example, consider an election year in which the economy is experiencing high inflation. Voters may blame the incumbent party, lowering their chances at reelection. Alternatively, voters may trust one of the political parties more (regardless of incumbency) to handle inflation. With the goal of estimating the affects of changes in election probabilities on stock markets, these possibilities introduce severe endogeneity concerns. Two potential approaches to mitigate this identification threat are (i) taking a narrative approach by selecting changes around events plausibly exogenous to the economy; and (ii) orthogonalizing the election probability movements to developments in the economy. Relative to the first approach, the second approach has the advantage of producing a less sparse series, and the drawback of introducing more noise. In this section I take the second approach, returning to the narrative approach in Section 4.2.
@@
where $\pi^{\text{R}}_{t}$ denotes the probability that the Republican candidate wins the next election as of day $t$, $\textbf{X}_{t}$ is a vector of contemporaneous economic series, and $\text{Pres}^{\text{R}}_{t}$ is dummy variable equal to 1 (0) when the current president is a Republican (Democrat). Note that by definition a symmetric linear equation exists when substituting $\pi^{\text{R}}_{t}$ for $\pi^{\text{D}}_{t}$, the probability of a Democratic candidate winning. A week (i.e., 5 business days) of autoregressive lags are included to purge the series of serial correlation. The remaining series represent potential sources of variation that may affect economic outcomes as well as election probabilities. For example, the set of interaction terms may affect election probabilities through the voters (dis)pleasure with the incumbent party, and affects financial and real conditions if each party reacts differently to economic conditions. If a sufficient set of predictors are included in $\textbf{X}_{t}$, then $\varepsilon_{t}$ identifies the variation in $\pi^{\text{R}}_{t}$ that comes from sources other than developments in the economy. Thus, it is important to carefully choose the content of this vector so that one may interpret $\varepsilon_{t}$ as an exogenous shock. In my baseline specification I set:
@@
where $\Delta i_{t}$ and $\Delta\text{sp}500_{t}$ are the 1-day percentage change in 2-year Treasury yields and the S&P 500 and $\Delta\text{emp}^{1m}_{t}$, $\Delta\text{cpi}^{1m}_{t}$, and $\Delta\text{ind}^{1m}_{t}$ give the most recent 1-month percentage change in employment, the CPI, and industrial production. $\mathbbm{1}_{e}(t)$, $\mathbbm{1}_{c}(t)$, and $\mathbbm{1}_{i}(t)$ are indicator functions equal to 1 on the day of the data release for each respective series. [^12] The employment situation from the preceding month is typically released on the first Friday of each month, while CPI and industrial production are released with roughly a 2-week lag. Thus, the last three components of $\textbf{X}_{t}$ take on values of 0 each day of the month except days where new information about the relevant macroeconomic series is released. It is critical to take the timing of these series’ release seriously since traditional stock markets (and plausibly election betting markets) react to news about the economy in real-time (Orphanides, 1992; Boyd et al., 2005; Lapp and Pearce, 2012). It is also important to account for data revisions to the initial releases to accurately reflect the information available at a given time. Therefore, for each component of $\textbf{X}_{t}$, I use values from the data vintages corresponding to each specific date, rather than relying on the most recent vintage, which introduces look-ahead bias. Note that equation (1) includes a week of lags for the base and interaction terms of $\textbf{X}_{t}$, which allows for a delayed market reaction. [^13] In unreported results I exclude the indicator functions so as to treat stale and new information the same. Results in Section 3 and 4 are unaffected. These components can be interpreted similarly to the “economic news” measure used by Bauer and Swanson (2023a, 2023b) to orthogonalize monetary policy shocks to information released between the Blue Chip Economic Indicators survey and FOMC announcements. One key difference between these two applications is the frequency of data – the monetary policy shocks are at the FOMC announcement frequency as opposed to the daily frequency.
@@
where $\hat{\varepsilon}_{t}$ represents the remaining 11.5% of variation in the data not explained by the model, which I interpret as an exogenous election surprise throughout the remainder of the paper.
@@ 3.3 Analysis of the Election Shocks
The first panel of Figure 5 plots $\text{Shock}_{t}$ over the entire sample. Positive shocks are associated with surprise increases to the Republicans’ probability of winning the upcoming election, while the converse is true for negative shocks. Shocks are set to equal 0 outside of election periods. The largest shocks are generally associated with election day, when $\pi^{\text{R}}_{t}$ goes to 0 or 1.151515There was still some uncertainty in the days after the 2000 and 2020 elections, but in my baseline specification I ignore this complication due to data limitations. The 2000 election poses a larger difficulty due to a month-long lawsuit over the outcome. In unreported results I exclude the 2000 election and the main results of this paper are largely unchanged. However, there is substantial heterogeneity in the size of these shocks. For instance the shocks associated with the Obama 2008 and Biden 2020 victories are each approximately -12pp, while the shock associated with the Trump 2016 victory is a staggering 71pp. In an informal sense, the 2008 and 2020 results were closer to a resolution of uncertainty than a “shock,” as Obama and Biden were widely expected to win. On the other hand, the 2016 result was a true surprise given the predictions of political analysts and media outlets. Also noteworthy, the magnitude of the shock associated with Trump’s 2024 victory is slightly overstated given the tightness of the race leading up to election day. This is a result of the original series measuring the odds of each candidate winning the popular vote, rather than the electoral college. [^16] In Section 4 this overstatement biases the magnitude of my results downward as the sizable stock market activity following the election are associated with a smaller surprise than the one suggested here. Overall, the shocks associated with election days realistically differentiate between the degree of surprise associated with each result. This feature of the series is critical, as stock markets are likely to react differently to expected and unexpected election outcomes.
@@ 4.1 Stock Markets
where $y^{i}_{t}$ is the log price of a sector stock index, $\Delta y^{i}_{t+h}$ is an $h$-day ahead long difference, $\Delta y^{i,1m}_{t}$ is the one-month change in $y^{i}_{t}$, and $\gamma^{i}_{h}$ are the coefficients of interest. Regressions are specified in differences to reduce bias in the estimates of $\gamma^{i}_{h}$ per the recommendations of Piger and Stockwell (2023) (also see Jordà and Taylor, 2024). For each sector and horizon I estimate equation (4) using weighted OLS, where the weight is given by the IEM trade volumes on day $t$ [^17] Trade volumes are plotted in Figure B2. Energy and defense prices are available for the entire sample, while clean energy prices are available starting in November 2003. Additionally, I exclude clean energy prices from September 2008 through December 2008 in the estimation due to a near 60% decrease during this period associated with the financial crisis. In unreported results I include these dates and point estimates are largely unchanged. In general, weighting by trade volumes reduces the variance of the estimates of $\gamma^{i}_{h}$, since election probability movements on days with low trade volumes may be driven by noise. In other words, it is advantageous to place more emphasis on observations with high trade volumes since these observations are associated with significant election events and dates closer to election day when markets are more tuned into election news. [^18] One drawback of this approach is that take place in non-election periods are excluded from the estimation. In Section 4.2 I explore an alternative approach that circumvent this weakness. An identification concern with examining far out responses is that changes in election probabilities are correlated with future political power, which in turn may result in the signing of legislation with direct affects on stock markets. To avoid this challenge I focus on the responses only up to 65 business days (3 months) after the shocks, which is approximately the distance between election day and inauguration.
Figure 6 plots the daily response of the energy, clean energy, and defense asset prices to a 10pp “Republican Shock,” i.e., an election shock that produces a contemporaneous 10pp increase to the the Republican candidate’s probability of winning the election. [^19] Figure D1 plots the response of $\pi^{\text{R}}_{t}$. Overall, election shocks have persistent effects on election probabilities. The black lines are point estimates and the shaded areas represent the 68% and 90% confidence intervals. Standard errors are estimated using the Newey-West HAC-consistent estimator with a bandwidth of $\frac{3}{4}\text{T}^{\frac{1}{3}}$, where T is the number of observations.
@@
Energy prices increase rapidly in response to the shock, peaking at about 1pp higher than baseline levels roughly four weeks after the shock. The effect of the shock is also persistent – at the furthest horizon prices remain significantly above baseline levels. Clean energy prices respond with a similar magnitude in the opposite direction, albeit with larger confidence intervals. Nonetheless, the response is significant in the short-term, and the point estimates remain well below baseline levels at the furthest horizon. The mirrored movements of energy and clean energy prices are reassuring – if the election shocks were confounded with, for instance, demand shocks, the two stocks would likely move in the same direction. Finally, defense prices display a similar pattern to energy prices peaking at about 1.5pp higher than baseline levels. Overall, these results substantiate the narrative evidence provided in Section 2 while providing additional insight about the medium-term affects of election news.
@@
A crude approach relative to the one taken here is to focus only on election outcomes by replacing Shockt with a variable equal to 1 (-1) on the day that a Republican (Democratic) candidate wins an election. Figure D2 plots the responses to this variable. The response of clean energy prices remains qualitatively similar, while the measure fails to capture persistent changes to defense and energy prices. This suggests that considering the entire election sample and the degree of “surprise” of each election outcome implied by the election probabilities is important. This is unsurprising given the anecdotal evidence provided in Figure 3 – asset prices on election day already reflect market expectations.
@@
The main results are robust to other sensible exercises. For example, a “one-step” alternative to constructing the election shocks is to estimate the response of asset prices to $\pi^{\text{R}}_{t}$ while controlling for the right-hand-side components of equation (1). Figure D3 plots the results under this specification, with similar results. [^20] By the Frisch-Waugh-Lovell theorem this approach is nearly analytically equivalent to my baseline specification. The two are not the same due to the inclusion of $\Delta y^{i,1m}_{t-1}$ in equation (4), the use of weights, and in the case of clean energy differences in the sample between the 1st and 2nd stages. In addition, two potential factors that feasibly confound the main results are irregularities in stock markets associated with the 2007-2008 financial crisis and the COVID-19 pandemic. The former of these events spanned the entire 2008 election cycle, while the latter spanned the majority of the 2020 election cycle. If rapid changes in asset prices related to one of these two events accompanied large swings in election probabilities, my estimates could be biased in either direction. To alleviate this concern I re-estimate equation (4) after dropping observations from the 2008 and 2020 elections and plot the resulting impulse responses in Figure D4. Reassuringly, the results are largely unchanged.

### v2

@@ 3.1 Data
The primary data used to construct the election shocks come from Iowa Electronic Markets (IEM), a futures contract market operated by the University of Iowa for research purposes. IEM publicly provides data on its winner take all futures betting markets for presidential races. Betting markets have been shown to more accurately predict elections than polling data (Berg, Nelson and Rietz, 2008). Relative to other betting markets, IEM has several desirable properties: (i) bets are capped at \$500, mitigating the influence of single traders; (ii) the markets have consistently high trading volumes, especially in the last few months of an election; (iii) the data are at a daily frequency, which is vital for studying stock markets; and (iv) market data is available for every presidential election since 1988888I focus on elections starting in 2000 due to the lack of availability of trading volumes and sectoral price indices for earlier elections., and in all election years data are available starting in June or earlier and extend through election day in November. One drawback of these data are that the payouts are based on which party wins the popular vote rather than the electoral college. In other words, the implied probability of a candidate winning based on these futures prices are biased if her chances of winning the popular vote are different from her chances of winning the election. Indeed, this is particularly relevant for recent elections, e.g., based on other betting markets Harris’s predicted probability of winning the popular vote was much higher than her probability of winning. [^9] For example, odds from Polymarket.com the day before the election implied Harris had a 71 percent chance to win the popular vote, but a 41 percent chance to win the presidency. While the last three elections had electoral college bias in favor of the Republican candidates, there is no clear historical systematic bias (Erikson, Sigman and Yao, 2020; Coleman and Kondik, 2023). With this in mind, to construct the election shocks I rely on the implicit correlation between innovations to both a candidates’ probability of winning the popular vote and the probability that they win electoral college.
@@
Figure 4 plots the 2024 Republican and Democratic candidates’ implied winning probabilities from June 1st through November 4th, the day before the election. The probabilities sum to 100, i.e., an increase to the Republicans’ probabilities are associated with a 1-to-1 decrease in the Democrats’ probabilities. Democratic candidates were each favorites to win the popular vote throughout the sample, although there is significant heterogeneity in the magnitude of this advantage. Vertical lines indicate notable events that are associated with substantial movements in the data. The first and most notable event is the aforementioned debate between Biden and Trump, which is associated with a contemporaneous 15pp increase to Trump’s probability of winning, albeit leveling off roughly 10pp higher than pre-debate levels. On July 11, Biden held a news conference that received mixed reactions, but overall was seen as a positive signal relative to his debate performance. This event is associated with a moderate increase to the Democrats’ odds of presidency. Just two days later, Trump survived an assassination attempt during a rally in Pennsylvania, which bolstered his election probabilities. A medium-term increase in the Democrats likelihoods appears to be linked to Biden dropping out of the race and endorsing Harris on July 21. Finally, the aforementioned debate between Harris and Trump is associated with a steady increase in the Democrats’ odds, which peak about 16pp higher roughly two weeks after the debate. See Figure B1 for the probabilities of other elections in the sample and Table LABEL:tabc2 for information on other identified events. Overall, these data appear to adequately track fluctuations in each party’s election probabilities around significant events.
@@ 3.2 Construction of the Election Shocks
The goal of this section is to construct a shock series that captures the types of fluctuations in election probabilities that are emphasized in Figure 4. However, a significant challenge to this exercise is the possibility that movements in election probabilities also reflect changes in voter preferences given changes to the state of financial markets or the real economy. For example, consider an election year in which the economy is experiencing high inflation. Voters may blame the incumbent party, lowering their chances at reelection. Alternatively, voters may trust one of the political parties more (regardless of incumbency) to handle inflation. With the goal of estimating the effects of changes in election probabilities on stock markets, these possibilities introduce severe endogeneity concerns. Two potential approaches to mitigate this identification threat are (i) taking a narrative approach by selecting changes around events plausibly exogenous to the economy; and (ii) orthogonalizing the election probability movements to developments in the economy. Relative to the first approach, the second approach has the advantage of producing a less sparse series, and the drawback of introducing more noise. In this section I take the second approach, returning to the narrative approach in Section 4.2.
@@
where $\pi^{\text{R}}_{t}$ denotes the probability that the Republican candidate wins the next election as of day $t$, $\textbf{X}_{t}$ is a vector of contemporaneous economic series, and $\text{Pres}^{\text{R}}_{t}$ is dummy variable equal to 1 (0) when the current president is a Republican (Democrat). A week (i.e., 5 business days) of autoregressive lags are included to purge the series of serial correlation. The remaining series represent potential sources of variation that may affect economic outcomes as well as election probabilities. For example, the set of interaction terms may affect election probabilities through the voters (dis)pleasure with the incumbent party, and affects financial and real conditions if each party reacts differently to economic conditions. If a sufficient set of predictors are included in $\textbf{X}_{t}$, then $\varepsilon_{t}$ identifies the variation in $\pi^{\text{R}}_{t}$ that comes from sources other than developments in the economy. Thus, it is important to carefully choose the content of this vector so that one may interpret $\varepsilon_{t}$ as an exogenous shock. In my baseline specification I set:
@@
where $\Delta i_{t}$ and $\Delta\text{sp}500_{t}$ are the 1-day percentage change in 2-year Treasury yields and the S&P 500 and $\Delta\text{emp}^{1m}_{t}$, $\Delta\text{cpi}^{1m}_{t}$, and $\Delta\text{ind}^{1m}_{t}$ give the most recent 1-month percentage change in employment, the CPI, and industrial production. $\mathbbm{1}_{e}(t)$, $\mathbbm{1}_{c}(t)$, and $\mathbbm{1}_{i}(t)$ are indicator functions equal to 1 on the day of the data release for each respective series. [^11] The employment situation from the preceding month is typically released on the first Friday of each month, while CPI and industrial production are released with roughly a 2-week lag. Thus, the last three components of $\textbf{X}_{t}$ take on values of 0 each day of the month except days where new information about the relevant macroeconomic series is released. It is critical to take the timing of these series’ release seriously since traditional stock markets (and plausibly election betting markets) react to news about the economy in real-time (Boyd, Hu and Jagannathan, 2005; Lapp and Pearce, 2012). It is also important to account for data revisions to the initial releases to accurately reflect the information available at a given time. Therefore, for each component of $\textbf{X}_{t}$, I use values from the data vintages corresponding to each specific date, rather than relying on the most recent vintage, which introduces look-ahead bias. Note that equation (1) includes a week of lags for the base and interaction terms of $\textbf{X}_{t}$, which allows for a delayed market reaction. [^12] In unreported results I exclude the indicator functions so as to treat stale and new information the same. Results in Sections 3 and 4 are unaffected. These components can be interpreted similarly to the “economic news” measure used by Bauer and Swanson (2023a, b) to orthogonalize monetary policy shocks to information released between the Blue Chip Economic Indicators survey and FOMC announcements. One key difference between these two applications is the frequency of data – the monetary policy shocks are at the FOMC announcement frequency as opposed to the daily frequency.
@@
where $\hat{\varepsilon}_{t}$ represents the remaining 11.5 percent of variation in the data not explained by the model, which I interpret as an exogenous election surprise throughout the remainder of the paper.
@@ 3.3 Analysis of the Election Shocks
The first panel of Figure 5 plots $\text{Shock}_{t}$ over the entire sample. Positive shocks are associated with surprise increases to the Republicans’ probability of winning the upcoming election, while the converse is true for negative shocks. Shocks are set to equal 0 outside of election periods. The largest shocks are generally associated with election day, when $\pi^{\text{R}}_{t}$ goes to 0 or 1.141414There was still some uncertainty in the days after the 2000 and 2020 elections, but in my baseline specification I ignore this complication due to data limitations. The 2000 election poses a larger difficulty due to a month-long lawsuit over the outcome. In unreported results I exclude the 2000 election and the main results of this paper are largely unchanged. However, there is substantial heterogeneity in the size of these shocks. For instance the shocks associated with the Obama 2008 and Biden 2020 victories are each approximately -12pp, while the shock associated with the Trump 2016 victory is a staggering 71pp. In an informal sense, the 2008 and 2020 results were closer to a resolution of uncertainty than a “shock,” as Obama and Biden were widely expected to win. On the other hand, the 2016 result was a true surprise given the predictions of political analysts and media outlets. Also noteworthy, the magnitude of the shock associated with Trump’s 2024 victory is slightly overstated given the tightness of the race leading up to election day. This is a result of the original series measuring the odds of each candidate winning the popular vote, rather than the electoral college. [^15] In Section 4 this overstatement biases the magnitude of my results downward as the sizable stock market activity following the election are associated with a smaller surprise than the one suggested here. Overall, the shocks associated with election days realistically differentiate between the degree of surprise associated with each result. This feature of the series is critical, as stock markets are likely to react differently to expected and unexpected election outcomes.
@@ 4.1 Stock Markets
where $y^{i}_{t}$ is the log price of a sector stock index, $\Delta y^{i}_{t+h}$ is an $h$-day ahead long difference, $\Delta y^{i,1m}_{t}$ is the one-month change in $y^{i}_{t}$, and $\gamma^{i}_{h}$ are the coefficients of interest. Regressions are specified in differences to reduce bias in the estimates of $\gamma^{i}_{h}$ per the recommendations of Piger and Stockwell (2025) (also see Jordà and Taylor, 2025). For each sector and horizon I estimate equation (4) using weighted OLS, where the weight is given by the IEM trade volumes on day $t$ [^16] Trade volumes are plotted in Figure B2. Energy and defense prices are available for the entire sample, while clean energy prices are available starting in November 2003. Additionally, I exclude clean energy prices from September 2008 through December 2008 in the estimation due to a near 60 percent decrease during this period associated with the financial crisis. In unreported results I include these dates and point estimates are largely unchanged. In general, weighting by trade volumes reduces the variance of the estimates of $\gamma^{i}_{h}$, since election probability movements on days with low trade volumes may be driven by noise. In other words, it is advantageous to place more emphasis on observations with high trade volumes since these observations are associated with significant election events and dates closer to election day when markets are more tuned into election news. [^17] One drawback of this approach is that take place in non-election periods are excluded from the estimation. In Section 4.2 I explore an alternative approach that circumvents this weakness. An identification concern with examining far out responses is that changes in election probabilities are correlated with future political power, which in turn may result in the signing of legislation with direct effects on stock markets. To avoid this challenge I focus on the responses only up to 65 business days (3 months) after the shocks, which is approximately the distance between election day and inauguration.
Figure 6 plots the daily response of the energy, clean energy, and defense asset prices to a 10pp “Republican Shock,” i.e., an election shock that produces a contemporaneous 10pp increase to the the Republican candidate’s probability of winning the election. [^18] Figure D1 plots the response of $\pi^{\text{R}}_{t}$. Overall, election shocks have persistent effects on election probabilities. The black lines are point estimates and the shaded areas represent the 68 and 90 percent confidence intervals. Standard errors are estimated using the Newey-West HAC-consistent estimator with a bandwidth of $\frac{3}{4}\text{T}^{\frac{1}{3}}$, where T is the number of observations.
@@
Energy prices increase rapidly in response to the shock, peaking at about 1 percent higher than baseline levels roughly four weeks after the shock. The effect of the shock is also persistent – at the furthest horizon prices remain significantly above baseline levels. Clean energy prices respond with a similar magnitude in the opposite direction, albeit with larger confidence intervals. Nonetheless, the response is significant in the short-term, and the point estimates remain well below baseline levels at the furthest horizon. The mirrored movements of energy and clean energy prices are reassuring – if the election shocks were confounded with, for instance, demand shocks, the two stocks would likely move in the same direction. Finally, defense prices display a similar pattern to energy prices peaking at about 1.5 percent higher than baseline levels. Overall, these results substantiate the narrative evidence provided in Section 2 while providing additional insight about the medium-term effects of election news.
@@
A crude approach relative to the one taken here is to focus only on election outcomes by replacing Shockt with a variable equal to 1 (-1) on the day that a Republican (Democratic) candidate wins an election, and 0 otherwise. Figure D2 plots the responses to this variable. The response of clean energy prices remains qualitatively similar, while the measure fails to capture persistent changes to defense and energy prices. This suggests that considering the entire election sample and the degree of “surprise” of each election outcome implied by the election probabilities is important. This is unsurprising given the anecdotal evidence provided in Figure 3 – asset prices on election day already reflect market expectations.
The main results are robust to other sensible specifications. In the main specification I treat the election shock series as observed, ignoring estimation uncertainty involved in its construction when computing standard errors. Figure D3 plots the estimation results of a “one-step” alternative that estimates the response of asset prices to $\pi^{\text{R}}_{t}$ while controlling directly for the right-hand-side components of equation (1). [^19] By the Frisch-Waugh-Lovell theorem, the point estimates of this approach are nearly analytically equivalent to the baseline specification. They are different due to the inclusion of $\Delta y^{i,1m}_{t-1}$ in equation (4), the use of weights, and in the case of clean energy differences in the sample between the 1st and 2nd stages. The point estimates and confidence intervals do not meaningfully differ from the baseline, indicating the generated regressor problem is not consequential.
@@
Two potential factors that feasibly confound the main results are irregularities in stock markets associated with the 2007-2008 financial crisis and the COVID-19 pandemic. The former of these events spanned the entire 2008 election cycle, while the latter spanned the majority of the 2020 election cycle. If rapid changes in asset prices related to one of these two events accompanied large swings in election probabilities, my estimates could be biased in either direction. To alleviate this concern I re-estimate equation (4) after dropping observations from the 2008 and 2020 elections and plot the resulting impulse responses in Figure D4. Reassuringly, the results are largely unchanged.